# Flow Matching Meets PDEs: A Unified Framework for Physics-Constrained Generation

### v2

@@ 3 Experimental Setup
We evaluate our method on three benchmark problems. Across all experiments, we employ a diffusion transformer (DiT) backbone architecture (Peebles and Xie, 2023), with minor modifications detailed in Appendix F. For completeness, we also provide a comparison to the UNet from previous work (Ronneberger et al., 2015; Bastek et al., 2025), shown in Figure 11. The problems differ in difficulty and were chosen to cover the three types of physics residuals (steady-state, transient, and analytic) as outlined above. A complete description of datasets is provided in Appendix B.
@@
We begin with the two-dimensional Darcy flow problem to compare with previous work Bastek et al. (2025). The Darcy equations which models steady-state fluid flow through a porous medium. The solution comprises pressure $p$ and permeability $K$. We use the corresponding public dataset (Bastek, 2024), which contains 10k training and 1k validation samples, each of size $64\times 64$. It is worth noting that, while common in literature (Zhu and Zabaras, 2018; Jacobsen et al., 2025), this dataset lacks conditioning inputs, making it less representative of real-world application scenarios. The residual directly corresponds to the governing PDE:
$\mathcal{R}=\boldsymbol{\nabla}\cdot(K\boldsymbol{\nabla}p)+f=0$
@@
The second benchmark is the two-dimensional Kolmogorov flow over a $128\times 128$ spatial domain with periodic boundary conditions. The dataset is generated using a spectral solver and consists of velocity field snapshots for various Reynolds numbers in the range [100; 500][100;\ 500]. The Reynolds number conditions the data generation, influencing the turbulence scales and flow complexity. The training set includes data for 32 Reynolds numbers with 1024 temporal snapshots each. The test set additionally includes 16 unseen Reynolds numbers. The data distribution reflects the temporal variation within each flow regime. The prediction target consists of the two velocity components, which are expected to satisfy the conservation of mass via:
$\mathcal{R}=\boldsymbol{\nabla}\cdot\boldsymbol{U}=0$
@@
The final and most complex setup involves spatio-temporal fields over a pitching NACA0012 airfoil, capturing the effects of dynamic stall in compressible flows. Each sample consists of $128\times 128$ distributions of absolute pressure, temperature, density, skin friction, and tangential velocity gradients. This physical model is highly relevant for real world cases such as the design and control of helicopter and wind turbine blades, where dynamic stall plays a critical role. Conditioning is performed on four parameters that define the operating conditions and airfoil motion. The training set comprises 128 base configurations, each perturbed 32 times to model uncertainty, while the validation set includes 16 unseen configurations. Physical consistency across fields is enforced through two analytical, point-wise residual constraints: $\mathcal{R}_{\text{ig}}$ imposes the ideal gas law, while $\mathcal{R}_{\tau}$ minimizes the skin friction starting from the tangential velocity gradients and the dynamic viscosity computed with Sutherland’s law.
$\mathcal{R}_{\text{ig}}$ $=P-\rho RT$
$\mathcal{R}_{\tau}$ $=\tau_{w}-\mu_{0}\frac{T_{0}+S}{T+S}\left(\frac{T}{T_{0}}\right)^{\frac{3}{2}}\sqrt{\left(\frac{\partial u}{\partial x}\right)_{n=0}^{2}+\left(\frac{\partial u}{\partial y}\right)_{n=0}^{2}}$
@@ 4 Results
The key findings of this study highlight the framework’s effectiveness in reducing physical residual errors while closely matching the data distribution. For the conditioned cases, the Kolmogorov flow and dynamic stall, we further assess the mean and standard deviation of each conditioned distribution.
@@ 4.1 Darcy flow
We benchmark our proposed framework, PBFM, against five alternative models: the FM baseline trained without incorporating physical constraints, the PBFM-1 model that includes the additional loss term but omits unrolling. Additionally, we compare with DM as DDPM baseline, the state-of-the-art PIDM-ME model Bastek et al. (2025), and the CoCoGen model of Jacobsen et al. (2025). The latter is an interesting variant, as it, contrary to the other algorithms, employs residual gradients during inference.
Figure 3: Darcy flow validation over 1024 samples using 20 FM steps. The left panel shows the residual MAE as a function of training steps, followed by visualizations of the residual error (error bars refer to min-max values within the validation dataset samples), pressure, and permeability distributions. Importantly, the low residual error of PIDM-ME is attributed to its partial coverage of pressure, and permeability, distributions (shown right). Both graphs highlight the high distributional accuracy of PBFM.
@@
We begin by evaluating overall performance, focusing on residual mean absolute error (MAE), along with pressure and permeability distributions. The results are summarized in Figure 3. The first panel illustrates how residual error evolves with the number of unrolling steps used during training. The bar plot shows that MAE decreases significantly from approximately 1.7 for FM to about 0.6 for PBFM. Notably, the improvement achieved by introducing the additional loss term is comparable in magnitude to the improvement gained via unrolling. When comparing across methods, we observe that the DM baseline performs similarly to FM. CoCoGen produces an $\approx\!10\%$ lower residual error than FM but ca. $20\%$ higher than PBFM-1, at the same time incurring a higher compute cost for inference. While PIDM-ME yields the lowest residual error, this performance comes at a steep cost: As shown in the last two panels, the residual scores for PIDM-ME are associated with a severely degraded pressure distribution, characterized by values constrained within the narrow range $\pm 0.2$. This artificially reduces the residual due to the lower magnitude of $P$, rather than improved fidelity. In comparison, all variants of our framework more faithfully reproduce the true pressure distribution. Regarding permeability, the DM baseline, as well as CoCoGen, exhibit a loss of distributional fidelity, particularly at the extrema. While our framework also shows a slight degradation for the least probable values, it successfully recovers the maximum $K$ values around 30, which are approximately six orders of magnitude less frequent than the central mode of the distribution. Figure 4 visualizes the trend of decreasing residual error as the FM model is progressively improved.
Figure 5: Darcy flow MAE residual (error bars refer to min-max values over validation set) and Wasserstein distance over FM steps for 1024 samples.
To complete the analysis on the Darcy dataset, we evaluated the PBFM-1 framework for sampling time $t$ during training with the logit-normal distribution, comparing it to the standard uniform distribution. We also analyze the effect of different samplers: deterministic and stochastic. Figure 5 reports the residual error along with the Wasserstein distance (Panaretos and Zemel, 2019; SciPy, ) as a function of the number of FM steps. The use of the logit-normal distribution does not appear to offer any particular advantage over the uniform distribution, in terms of either residual error or distribution. In contrast, switching to a stochastic sampler reduces the residual error if the number of function evaluations exceeds 10. However, this improvement diminishes when the number of FM steps becomes too large, as the distributional quality worsens. Particularly, with 20 integration steps, the proposed sampler reduces the residual error by a significant factor of $3\times$, while the Wasserstein distance is only 30% lower for the deterministic sampler. Further results showing pressure and permeability samples and an additional UNet are reported in Figure 10 and 11 of the appendix.
@@ 4.2 Kolmogorov flow
Kolmogorov flow differs from the Darcy case as it is conditioned on the Reynolds number, and in this case the physical residual only matches parts of the underlying PDE. Here, the PBFM model reduces the average residual MSE by a factor of $1.68\times$ compared to the baseline, while the PBFM-1 approach without unrolling offers only limited improvement of $1.05\times$, as shown Figure 6. This difference is clearly reflected in the panel on the left, which visually depict the residual MSE. When examining the velocity distributions, all FM variants yield nearly overlapping profiles, although none fully capture the extrema of the reference data. In the appendix, Figure 13 and 14 report the instantaneous velocity fields along the divergence and an example of mean and standard deviation, respectively. This test case highlights the importance of accurate samples via unrolling to fully leverage the physics residuals at training time.
Figure 6: Kolmogorov flow residual and distribution comparison over 128 samples using 20 FM steps. Examples of residual MSE for the three analyzed approaches. Violin plot distributions refer to the different conditioning values within the validation dataset.
@@ 4.3 Dynamic Stall
The dynamic stall case presents additional challenges, including two physical constraints of increased complexity and six predicted fields. The underlying phenomena are highly non-linear and governed by a system of four time-dependent PDEs.
Figure 7 summarizes the key results, including the residual MSE, Wasserstein distance. Focusing first on the residual error, we observe a consistent improvement when progressing from the FM baseline to PBFM, with the PBFM-1 providing a $2\times$ reduction compared to the baseline. The incorporation of residual loss and unrolling not only reduces the average error of a $8\times$ factor but also sharpens the residual error distribution, concentrating values closer to the mean. In terms of data distribution, the Wasserstein distance reveals that while the PBFM-1 framework alone may degrade distributional accuracy, the addition of unrolling effectively regularizes the predictions, narrowing the gap with the ground truth distribution.
Figure 7: Comparison of the proposed approaches of physical residual MSE, Wasserstein distance, for dynamic stall. Each condition includes 128 samples with 20 FM steps. Violin plots show distributions over the validation dataset.
It is worth noting that MSEs of the predicted mean and standard deviation are on par for all methods (details are given in the appendix), indicating that the improvements of PBFM in residual and distributional accuracy are achieved without compromising first- and second-order statistics.
Figure 8 presents a visual comparison: the left panel illustrates the mean and standard deviation of the pressure field $P$ and wall shear stress $\tau_{w}$, while the right panel shows the residual error magnitude. PBFM accurately captures both the average behavior and the associated fluctuations, maintaining consistency with the minimum and maximum bounds of each variable. Regarding residuals, the PBFM-1 model demonstrates a reduction in the average error relative to the baseline. However, it is worth noting that the high-error regions do not show significant improvement. In contrast, incorporating the unrolling strategy leads to a notable reduction even in zones where the error is largest.
@@
Figure 8: Dynamic stall example comparison of mean and standard deviation for PBFM framework computed over 128 samples using 20 FM steps. Examples of residual MSE for the three analyzed approaches.
Performance
At training time, evaluating physical residuals and unrolling both lead to slight increases in computational load. Importantly, the increased accuracy of PBFM does not negatively affect performance at inference time. While reliably comparing computational performance is difficult, the dynamic stall scenario roughly indicates benefits of the proposed generative model: the CPU-based data generating solver takes 76 minutes on average for a single simulation, while the trained PBFM model requires only 4 minutes to generate 128 samples on the same CPU architecture and 0.2 seconds on an A100 GPU. Detailed measurements are provided in Table 4.
5 Related work
Numerous deep learning methods have been developed to address complex problems in physics and engineering (Morton et al., 2018; Wang et al., 2020; Sanchez-Gonzalez et al., 2020; Thuerey et al., 2020). More recently, there has been growing interest in bringing the foundation model paradigm to scientific machine learning. Efforts such as PDEformer (Ye et al., 2024), Poseidon (Herde et al., 2024), Aurora (Bodnar et al., 2024), and Unisolver (Zhou et al., 2025) aim to build models with broad generalization capabilities across diverse physical systems. In parallel, specialized transformer architectures tailored for PDEs have also been proposed, including OFormer (Li et al., 2023), Transolver (Wu et al., 2024), Fengbo (Pepe et al., 2025), and PDE-Trasformer (Holzschuh et al., 2025).
Focusing on diffusion models in the physics and engineering fields, representative applications include the generation and design of new molecules and drugs (Guo et al., 2024; Schneuing et al., 2024; Bose et al., 2024), the simulation of particle trajectories in collider experiments (Mikuni et al., 2023), solving inverse PDE-problems (Holzschuh et al., 2023), and the modeling of particle motion in turbulent flows (Li et al., 2024). Alongside these efforts, several works have sought to embed physical knowledge or constraints directly into the diffusion process to enhance model performance. For instance, Huang et al. (2024) presented a framework to solve PDEs from partial observations by filling in missing information using generative priors. Rixner and Koutsourelakis (2021) formulated a probabilistic generative model enforcing physical constraints through virtual observables.
More aligned with our work, Shu et al. (2023) proposed a framework that conditions the diffusion process on the residual gradient both at training and inference time, with a focus on super-resolution and reconstruction tasks based on random measures of turbulent flows. Despite being one of the first work in this direction, they do not directly enforce the residual. Similarly, Jacobsen et al. (2025) introduced CoCoGen, a method that employs the governing PDEs during inference, while leaving the training procedure unchanged. The approach improves the physical residual but slow down inference which is crucial in real applications. Giannone et al. (2023) demonstrated that aligning the sampling trajectory of diffusion models with trajectories obtained from physics-based iterative optimization methods can improve the quality of generated designs; however, their approach is limited to topology optimization problems rather than general PDEs. Another topology optimization work that enforces physical constraint is proposed by Mazé and Ahmed (2023). They included in the sampling process a surrogate neural network predicting and enforcing the compliance of the topology under given constraints. Finally, Bastek et al. (2025) proposed physics-based diffusion models (PIDM), which incorporate an additional term during diffusion training to minimize physical residuals. Despite their formal derivation of the objective, no concrete solution is proposed to address the inherent conflict between the generative and physical objectives.
6 Conclusions
[…]

### v3

@@ 4 Experimental Setup
We evaluate the generative performance of our method, denoted with Physics Based Flow Matching (PBFM) in the following, on three benchmark problems. Across all experiments, we employ a diffusion transformer (DiT) backbone architecture (Peebles & Xie, 2023), with minor modifications detailed in Appendix G. For completeness, we also provide a comparison to the UNet from previous work (Ronneberger et al., 2015; Bastek et al., 2025), shown in Fig. 10. Although our method is agnostic to the type of physical residual, we organize the three benchmarks according to the residual categories described above: steady-state, transient, and analytic. For each case, the residual is computed using: finite differences for Darcy flow, FFT-based methods for Kolmogorov flow, and point-wise evaluation for dynamic stall. A complete description of datasets is provided in Appendix C.
@@
We begin with the two-dimensional Darcy flow problem. The Darcy equation which models steady-state fluid flow through a porous medium. The solution comprises pressure $p$ and permeability $K$. We use the corresponding public dataset (Bastek, 2024), which contains 10k training and 1k validation samples, each of size $64\times 64$. It is worth noting that, while common in literature (Zhu & Zabaras, 2018; Jacobsen et al., 2025), this dataset lacks conditioning inputs, making it less representative of real-world application scenarios. The residual directly corresponds to the governing PDE: $\mathcal{R}=\boldsymbol{\nabla}\cdot(K\boldsymbol{\nabla}p)+f=0$
@@
The second benchmark is the two-dimensional Kolmogorov flow over a $128\times 128$ spatial domain with periodic boundary conditions. The dataset is generated using a spectral solver and consists of velocity field snapshots for various Reynolds numbers in the range [100; 500][100;\ 500]. The Reynolds number conditions the data generation, influencing the turbulence scales and flow complexity. The training set includes data for 32 Reynolds numbers with 1024 temporal snapshots each. The test set additionally includes 16 unseen Reynolds numbers. The data distribution reflects the temporal variation within each flow regime. The prediction target consists of the two velocity components, which are expected to satisfy the conservation of mass via: $\mathcal{R}=\boldsymbol{\nabla}\cdot\boldsymbol{U}=0$
@@
The final and most complex setup involves spatio-temporal fields over a pitching NACA0012 airfoil. This case captures the effects of dynamic stall, a complex, unsteady phenomenon in aerodynamics causing large-scale fluctuations and loads. As such, this physical model is highly relevant for real world cases such as helicopter and wind turbine blades, where dynamic stall plays a critical role. The solutions are obtained solving the compressible Navier-Stokes equations and the flow fields present shock waves and detailed flow phenomena that are challenging to capture. Each sample consists of $128\times 128$ distributions of six quantities: absolute pressure, temperature, density, skin friction, and tangential velocity gradients (x and y). Conditioning is performed on four parameters that define the operating conditions and airfoil motion. The training set comprises 128 base configurations, each perturbed 32 times to model uncertainty, while the validation set includes 16 unseen configurations. Physical consistency across fields is enforced through two analytical, point-wise residual constraints: $\mathcal{R}_{\text{ig}}$ imposes the ideal gas law, while $\mathcal{R}_{\tau}$ minimizes the skin friction with Sutherland’s law.
$\mathcal{R}_{\text{ig}}=P-\rho RT\ ,\ \ \ \ \ \ \ \ \mathcal{R}_{\tau}=\tau_{w}-\mu_{0}\frac{T_{0}+S}{T+S}\left(\frac{T}{T_{0}}\right)^{\frac{3}{2}}\sqrt{\left(\frac{\partial u}{\partial x}\right)_{n=0}^{2}+\left(\frac{\partial u}{\partial y}\right)_{n=0}^{2}}$ (5)
@@ 5 Results
We evaluate our proposed framework, PBFM (Baldan et al., 2025a), in direct comparison with six representative baselines: flow matching with optimal transport FM-OT trained without physical constraints, PIDM-ME (Bastek et al., 2025), CoCoGen (Jacobsen et al., 2025), DiffusionPDE (Huang et al., 2024), D-Flow (Ben-Hamu et al., 2024), and ECI (Cheng et al., 2025). ECI only enforces BCs but not the PDE due to the limitations of the method. For each method, we report physical residual error, distributional fidelity via Wasserstein distance and Jensen-Shannon divergence, number of function evaluations, and inference wall-clock time per sample. For the conditional benchmarks, Kolmogorov flow and dynamic stall, we additionally evaluate the mean and standard deviation of the predicted distributions for each conditioning input. Comprehensive sample visualizations and further analysis are provided in Appendix D.
Figure 3: Pareto front of physical residual MAE vs. Wasserstein distance for Darcy flow for the proposed methods.
@@ 5.1 Darcy flow
Unlike previous works that report only individual metrics, we present a comprehensive evaluation of Pareto optimality in terms of physical residual error and distributional accuracy. The latter is quantified by the Wasserstein distance (WD) (Villani, 2008; SciPy, 2025). Fig. 3 visualizes the trade-off for all compared methods. The FM-OT baseline yields a WD of 0.059 but incurs a high residual error of 4.159. DiffusionPDE achieves similar WD with marginally improved residuals. CoCoGen and D-Flow are strictly dominated by our approach, exhibiting both higher residuals and inferior distributional metrics. PIDM-ME attains the lowest residual (0.022), which comes at the expense of a substantially degraded distribution (WD 3.103), reflecting poor generative fidelity. In contrast, PBFM achieves a residual of 0.838 and WD of 0.138, demonstrating a more favorable balance between physical and generative objectives. A visual representation of the residual errors is reported in Fig. 5. The previous methods align along a negative-slope trend, highlighting the intrinsic trade-off between residual minimization and distributional accuracy. PBFM advances the Pareto front, reducing residuals without a considerable increase in WD. Comprehensive metrics are reported in Table 2, including Jensen-Shannon divergence (JS) (Lin, 1991) as a distributional metric in addition to WD, the number of function evaluations (NFE), and inference time (IT). PBFM maintains competitive inference speed (0.101 s), similarly to FM-OT, as the only difference is the use of the stochastic sampler.
Table 2: Generative performance metrics for Darcy flow problem over 1024 samples. RE: physical Residual MSE, WD: Wassserstein Distance, JS: Jensen-Shannon divergence, NFE: Number of function evaluations, IT: Inference wall-clock time. Best and second best results.
Metric PBFM FM-OT CoCoGen PIDM DiffusionPDE D-Flow ECI§
RE 0.838 4.159 1.320 0.022 3.388 2.286 3.045
WD $\cdot 10^{2}$ 0.138 0.059 0.249 3.103 0.089 0.147 2.892
JS $\cdot 10^{1}$ 0.256 0.131 0.360 3.179 0.139 0.237 2.818
NFE 20 20 100 100 20 20 20
IT $[s]$ 0.101 0.100 7.395 2.050 0.590 3.126 0.122
† This method uses the UNet architecture from Bastek et al. (2025).
‡ D-Flow method is unstable, samples are filtered using RE <5<5 condition resulting in 888 valid ones.
§ The physical constraint is applied only to the BCs ($\text{RE}_{\text{BC}}\approx 0$) but cannot be applied to the non-linear PDE.
Figure 4: Darcy flow validation over 1024 samples using 20 FM steps. Residual MAE as a function of training steps (left), residual MAE as a function of FM steps (right).
The first panel of Fig. 4 illustrates two aspects. On the one hand, it shows how residual error evolves with the number of unrolling steps used during training and it demonstrates that unrolling effectively mitigates Jensen’s gap, leading to lower residual errors. On the other hand, it highlights that only applying the ConFIG method during training offers a limited improvement. To further analyze the impact of unrolling, the second panel of Fig. 4 presents the residual error as a function of the number of FM steps used during inference. The results confirm that increasing the number of FM steps consistently reduces the residual error across all methods. Notably, PBFM delivers the lowest residual error and it is less sensitive to the number of FM steps, indicating that it achieves high physical accuracy even with fewer function evaluations. This is particularly advantageous in practical applications where computational resources are limited.
Table 3: Performance of stochastic sampler for Darcy flow problem over 1024 samples using 20 FM steps using PBFM. Noise resampling is enabled for $t<t^{*}$, details in Algorithm 2. $t^{*}=0.0$ coincides with deterministic sampler. Best and second best results.
Metric $t^{*}$
0.0 0.2 0.4 0.6 0.8 1.0
RE 0.828 0.838 0.970 0.876 0.774 0.632
WD $\cdot 10^{2}$ 1.470 0.138 0.150 0.187 0.302 0.316
JS $\cdot 10^{1}$ 0.919 0.256 0.257 0.313 0.361 0.379
To complete the analysis, we evaluate the impact of the stochastic sampler introduced in Algorithm 2. Table 3 reports the performance of PBFM with different thresholds $t^{*}$ for noise resampling during inference. Setting $t^{*}=0.0$ corresponds to the deterministic sampler, while $t^{*}=1.0$ implies resampling noise at every step. The results indicate that using the deterministic sampler yields comparable residual error to the stochastic approach, but with significantly worse distributional metrics. Focusing on values greater than 0, we observe that increasing $t^{*}$ leads to improved residual error, with the best performance at $t^{*}=1.0$. However, this comes at the cost of degraded distributional accuracy, as both WD and JS increase. A balanced choice of $t^{*}=0.2$ offers a good compromise, achieving low residuals while maintaining strong generative fidelity. Further results showing pressure and permeability samples and an additional UNet are reported in Fig. 9 and 10 of the appendix.
@@ 5.2 Kolmogorov flow
Kolmogorov flow presents a conditional generative modeling challenge, where the Reynolds number serves as the conditioning input and the physical residual is defined by the divergence of the velocity field, enforcing conservation of mass. Table 4 details the quantitative results. PBFM achieves the lowest physical residual error (1.362), outperforming all baselines. It also yields the best distributional metrics, with a low Wasserstein distance and Jensen-Shannon divergence, demonstrating excellent generative fidelity. The standard deviation error is slightly higher than FM-OT, but the overall distributional accuracy remains superior. Inference times for PBFM (98.97 ms) are competitive with FM-OT, while DiffusionPDE and D-Flow incur substantially higher computational costs due to FFT-based residual evaluation. D-Flow is unstable for Kolmogorov flow and does not produce valid samples despite the high computational cost.
Table 4: Generative performance metrics for Kolmogorov flow and dynamic stall problems using 20 FM steps. RE: physical Residual MSE, WD: Wasserstein Distance, JS: Jensen-Shannon divergence, MMSE: MSE of the mean fields, SMSE: MSE of the standard deviation fields, NFE: Number of function evaluations, IT: Inference wall-clock time. Best and second best results.
Dataset Metric PBFM OT-FM DiffusionPDE D-Flow PCFM
RE $\cdot 10^{1}$ 1.362 2.314 1.930 -
WD $\cdot 10^{1}$ 1.222 2.124 3.698 -
Kolmogorov JS $\cdot 10^{2}$ 7.440 12.49 19.39 -
Flow MMSE $\cdot 10^{4}$ 4.455 4.188 5.669 -
SMSE $\cdot 10^{4}$ 2.484 2.574 21.82 -
IT $[ms]$ 98.97 98.75 267.8 6431
RE $\cdot 10^{6}$ 0.339 11.02 12.20 11.32 0.143
WD $\cdot 10^{4}$ 1.814 2.707 2.509 3.484 4.013
Dynamic JS $\cdot 10^{2}$ 0.680 0.983 1.029 1.014 1.206
Stall MMSE $\cdot 10^{5}$ 1.490 2.791 2.626 2.507 5.669
SMSE $\cdot 10^{5}$ 0.874 1.458 1.236 1.372 7.674
IT $[ms]$ 60.47 59.75 171.7 138.9 3906
‡ D-Flow method is unstable for Kolmogorov Flow.
Qualitative results are provided in Appendix Fig. 12 and 13, which show representative instantaneous velocity fields, divergence residuals, and the predicted mean and standard deviation across the test set.
@@ 5.3 Dynamic Stall
Table 5: Metric comparison for different values of $\sigma_{\min}$ on the dynamic stall problem, 20 FM steps. Best and second best results.
Metric $\sigma_{\min}$
0.0 $\mathbf{10^{-4}}$ $\mathbf{10^{-3}}$ $\mathbf{10^{-2}}$
RE $\cdot 10^{6}$ 0.339 0.468 0.473 0.466
WD $\cdot 10^{4}$ 1.814 3.547 3.246 3.566
JS $\cdot 10^{2}$ 0.680 0.716 0.718 0.728
MMSE $\cdot 10^{5}$ 1.490 1.360 1.298 1.200
SMSE $\cdot 10^{5}$ 0.874 0.831 0.789 0.916
The dynamic stall case presents additional challenges, including two physical constraints of increased complexity and six predicted fields. The underlying phenomena are highly non-linear containing shock waves. Fig. 6 presents an example of predicted fields. Especially regions like the center exhibit strongly varying, small-scale shock waves that stochastically oscillate for perturbed operating conditions - a phenomenon that is important to capture for industrial applications (Baldan et al., 2025b).
Table 4 summarizes the quantitative results. PBFM outperforms all baselines across nearly all metrics, achieving the lowest distributional errors and MSE for both the mean and standard-deviation fields, which indicates that PBFM effectively captures both first- and second-order statistics of the complex flow fields. PCFM achieves the lowest residual error, but this comes at the cost of worse distributional metrics and higher mean and standard-deviation errors. Notably, PCFM incurs a significantly higher inference time ($\approx 65\times$ that of PBFM) due to the iterative correction process.
o Refer to caption
@@ 0.0 $\mathbf{10^{-4}}$ $\mathbf{10^{-3}}$ $\mathbf{10^{-2}}$
Figure 6: Dynamic stall, comparison of mean for PBFM framework computed over 128 samples using 20 FM steps.
Inference times for PBFM are comparable to FM-OT. DiffusionPDE and D-Flow incur higher computational costs because evaluating their physical residuals is more complex, although the cost difference is limited since the residuals are pointwise.
To investigate the impact of different values of $\sigma_{\min}$ on the dynamic stall problem, which present the lowest residual errors and is more sensible to this phenomenon, Table 5 analyzes the results. Setting $\sigma_{\min}=0$ yields the best distributional metrics and residual error underlines the importance of minimizing noise perturbation in this complex scenario. Increasing $\sigma_{\min}$ to $10^{-4}$ slightly degrades the metrics, while further increases to $10^{-3}$ and $10^{-2}$ lead to more pronounced declines in performance. This trend highlights the sensitivity of physical residuals to noise levels, emphasizing the need for careful tuning of $\sigma_{\min}$ in physics-constrained generative modeling. Nevertheless, with higher noise levels the model provides lower mean and standard deviation errors. A practical guideline is that adding Gaussian noise of scale $\sigma_{\text{min}}$ induces a residual MSE $\approx\sigma_{\text{min}}^{2}$ in a perfect reconstruction setting. For dynamic stall, this directly suggests an upper bound: $\sigma_{\text{min}}^{2}\lesssim 10^{-7}\Rightarrow\sigma_{\text{min}}\lesssim 3\times 10^{-4}$. In contrast, Kolmogorov and Darcy benchmarks have more relaxed residual requirements. Importantly, this sensitivity is not unique to our method; it is general to flow matching formulations, as the noise level determines the scale of the denoising target and hence the attainable physical accuracy.
Performance evaluation
We provide an overview of the training performance, focusing on the impact of the residual loss together with ConFIG, and the effect of unrolling. Table 6 reports the wall-clock time, in seconds, for a single training iteration on an NVIDIA A100 64 GB GPU. To reduce training computation time and memory footprint, backpropagation can be computed only through the final unrolling step; this does not significantly affect model performance. In Appendix J we provide a detailed comparison of this variant against full backpropagation through all unrolling steps.
Table 6: Comparison of wall-clock time in seconds for one training iteration and memory usage in GB on an NVIDIA A100 64GB GPU for the proposed approaches. Batch size is 64 for all cases. Inference time is unchanged.
FM PBFM PBFM PBFM PBFM
1 step 2 steps 3 steps 4 steps
Darcy [s] $4.28\cdot 10^{-2}$ $6.69\cdot 10^{-2}$ $9.78\cdot 10^{-2}$ $1.28\cdot 10^{-1}$ $1.59\cdot 10^{-1}$
[GB] 12.35 12.62 23.26 33.70 44.15
Kolmogorov [s] $1.13\cdot 10^{-1}$ $1.94\cdot 10^{-1}$ $3.02\cdot 10^{-1}$ $4.10\cdot 10^{-1}$ $5.18\cdot 10^{-1}$
[GB] 4.08 4.52 6.91 9.56 12.18
Dynamic [s] $4.29\cdot 10^{-2}$ $8.14\cdot 10^{-2}$ $1.18\cdot 10^{-1}$ $1.55\cdot 10^{-1}$ $1.90\cdot 10^{-1}$
[GB] 4.25 4.41 7.18 9.90 12.58
6 Discussion and conclusions
Limitations
PBFM delivers substantial improvements in physical and distributional accuracy, but its main limitation is increased computational and memory cost during training. Table 6 details the wall-clock time and memory usage for different unrolling steps. Incorporating the residual loss and ConFIG requires an additional backward pass, increasing training time and memory usage. Unrolling over 4 steps further increases memory consumption (up to $3\times$) and training time (up to $2.5\times$). However, these overheads are restricted to training; the inference speed of the method remains unaffected.
Discussion
PBFM is designed to enable rapid prediction of complex physical systems, particularly in scenarios where conventional solvers are prohibitively slow. Table 10 presents inference times for the dynamic stall benchmark, already employed for actual helicopter blade simulations. While a single simulation using a CPU-based solver requires approximately 76 minutes, the trained PBFM model generates 128 samples in just 4 minutes on CPU and only 0.2 seconds on GPU. This substantial acceleration demonstrates the practical utility of PBFM for large-scale uncertainty quantification and surrogate modeling tasks in scientific and engineering workflows. Without the proposed modifications, a trained model would not be sufficiently accurate for real-world applications.
Summary